evening shift summary 6/5/99

AG (ag@fnal.gov)
Sun, 06 Jun 1999 00:10:16 -0500 (CDT)

On Shift: Theo, Sasha

Hours of beam delivered at all intensities: 0.
Hours of beam delivered at nominal intensity: 0.
Hours of E832 data (special runs not included): 0.
Hours of special runs (laser scans, muon runs...): ~ 6.

Accounting of lost time: Switching back to E832 mode.

========================================================================
Runs ended this shift:

Run 13382 - PED
Run 13383 - LONG LASER SCAN
Run 13384 - COSMIC
Run 13388 - COSMIC in progress

Tapes brought to FCC this shift: 0

========================================================================

Accesses during shift: Open access during all shift.

========================================================================
Description of Shift:

15:30 Etot tripped of - Sydney is paged.

16:00 Ed changed HA1 and HA2 thresholds to 870mV, also x2 attenuator added
to HA2 to go back to 832 configuration

16:20 Ed and Hogan added spare fiber to CsI channel 1957.

16:30 PED run 13382 to update thresholds followed by

17:00 LONG LASER SCAN run 13383 - channel 1800 has large ped rms
( ~4 counts )

17:30 Sydney brings Etot back on. The trip occured due to TW crate
temperature of 92F, the high limit has been changed to 95F.

17:50 The laser scan is finished and COSMIC run 13384 started.

18:00 Channel 1957 with added fiber now has ~x2 too much light !
There are other channels that show larger gain drift after the work
on CsI ( although none of them were worked on! ):

1642 - 5 %
2733 - 3.9 %
2847 - 8 %

18:01 Regenerator is in and the mover is on.

18:20 "3X Kumquat Problem" - fixed by reloding kumquats with map-gen.

19:00 2 new ar/ethane botles connected ( east side )
1 new co2 bottle connected
DC2 exhaust line connected to main exhaust (flow is 0.7 cfh to DC2)

20:20 "kumquat problem" again - the same solution.

20:40 no triggers - DYC 53 FULL, trying to resolve the problem
kumquat 3X problem again, this time had to reset FSCC's...
still doesn't work.

Paged Jason L. - problem was resolved by resetting the KQ and
cycling power of the repeater downstairs by the Epicure window.

22:00 Ed requested to set NM2ABS2 ( common Be absorber )
to position 1 corresponding to 20.6" of the material.

22:30 The humidity went up due to the fact that after TRD's were moved
out they created a closed region next to the wall where the
humidity probe sits. The probe has been moved to a different place.
Now it looks fine.

23:30 Completed gas check list.